OpenStudy (anonymous):

Write the equation of the line passing through the point (1, 1/2) and perpendicular to the line y+4x=6

OpenStudy (mathstudent55):

Perpendicular lines have slopes that are negative reciprocals. That means that when you multiply the slopes together, you get -1. First you need to find the slope of y + 4x = 6. Then you find the slope of the perpendicular. To find the slope of y + 4x = 6, just solve that equation for y. Then you'll have the equation in the form y = mx + b, where m = the slope. y + 4x = 6 y = -4x + 6 m = -4. Slope of this line is -4; slope of the perpendicular is 1/4 since (-4)(1/4) = -1 Now you need to find the equation of the line that passes through point (1, 1/2) and has slope 1/4. Use the form y = mx + b, insert the point (1, 1/2) in for x and y, insert 1/4 for m and solve for b: y = mx + b 1/2 = (1/4)(1) + b 2 = 1 + b b = 1, so y = (1/4)x + 1